Sol Summary – December 7th

Sol: 10
Summary Title: Snow Day
Author’s name: Matt Eby
Mission Status: In-Simulation
Sol Activity Summary: Brief excursion to flip and reset the mirror experiment, drafting final report, adjusting plant supports in the GreenHab.
Look Ahead Plan: Balloon flight, cleaning, preparation for STEM outreach event.
Anomalies in work: None.
Weather: Calm winds throughout. Snow early, transitioning to drizzle for much of the day.
Crew Physical Status: Crew is in good shape, nothing of concern to note.
EVA: EVA# 1 led by the Crew Engineer, with Health and Safety Officer, traveled by foot to the mirror experiment. 1 hour duration.
Reports to be filed: Journalist, Commander, Operations, Photos, EVA report, EVA request, GreenHab
Support Requested: N/A

EVA Report – December 7th

EVA # 14
Author: B. Braun
Purpose of EVA: Reconfigure mirror coating experiment
Start time: 9:17 AM
End time: 9:58 AM
Narrative: EVA crew departed the airlock and proceeded on foot via the Hab driveway and Cow Dung Road to the area of the mirror location. Due to the snow, conditions were muddy and slippery along the road and across the terrain to the mirror experiment location. Once the team reached the mirror location, they unscrewed each of the mirror samples and flipped them over as directed by the experiment PI. After documenting the new setup with photos, the crew proceeded carefully back across the terrain to Cow Dung Road, then down the driveway to the Hab, where they paused to clear mud from their boots before entering the airlock.

Destination: Hab driveway to Cow Dung Road to mirror experiment location (near Cow Dung Road and Marble Ritual)
Coordinates (use UTM WGS 84): 4251000, 518500
Participants: Crew Engineer (Kowalski), HSO (Braun)
Road(s) and routes per MDRS Map: Cow Dung Rd (0110)
Mode of travel: Walking

Commander Report – December 7th

Sol: 10
Summary Title: Snow Day and EVA #14
Author’s name: Dr. Kristine Ferrone, Commander
Mission Status: Nominal
Commander’s Report:

Crew 269 awoke to a winter wonderland landscape this morning and snowy conditions. The crew enjoyed a quiche bake for breakfast and conducted a morning meeting to discuss how the weather would affect the day’s EVA plans. After a consultation with Tier 1 mission support, the crew decided to shorten the morning EVA #14 and continue to send 2 crew to perform repair operations on the mirror coating experiment payload. HSO and Engineer traversed on foot through the muddy terrain to repair the mirror payload and return swiftly to the Hab, while Green Hab Officer lead IVA comms. During the EVA, CDR and XO discussed end of mission reporting requirements and began to create content for those reports while Technology Officer continued to work data processing on the Aerospace computers. In the afternoon, HSO collected health and safety stats, crew completed exercise protocols, and Green Hab Officer tended to the plants in the Green Hab. Power was again limited due to weather conditions and the crew followed Tier 1 support instructions to conserve power. For dinner, the crew enjoyed black bean burgers on homemade sourdough buns with microgreens from the Green Hab.

Operations Report – December 7th

SOL: 10

Name of person filing report: Ashley Kowalski

Non-nominal systems: Toilet odor has returned

Notes on non-nominal systems: The odor from our toilet has returned. The water level in the toilet is getting high, so it needs to be emptied. However, since we had previous toilet issues, I would like to hear from Mission Support regarding the best plan of action. Should I empty the toilet system per normal guidelines, or do we need to do another full flushing out of the system as was performed two days ago? Please advise. We would like to perform this procedure tonight, but we will await your response before proceeding.


Spirit rover used: No
Hours: N/A
Beginning charge: N/A
Ending charge: N/A
Currently charging: N/A

Opportunity rover used: No.
Hours: N/A
Beginning charge: N/A
Ending charge: N/A
Currently charging: N/A

Curiosity rover used: No
Hours: N/A
Beginning charge: N/A
Ending charge: N/A
Currently charging: N/A

Perseverance rover used: No
Hours: N/A
Beginning charge: N/A
Ending charge: N/A
Currently charging: N/A

General notes on rovers: No additional notes/comments.

Summary of Hab operations:
WATER USE: ~375 gallons used to-date
Water (static tank): ~175 gallons (This is my best guess without a water gauge, but I am requesting Mission Support to let us know a more accurate value, if possible.)
Static tank pipe heater (on or off): On
Static tank heater (On or off): On
Toilet tank emptied: No.

Summary of internet: No issues with the internet. Connection is good.

Summary of suits and radios: We tested out the GreenHab Officer’s headset (headset #3) again today. It seemed to be working intermittently today. Please advise if there is anything you would like for us to test with the headset.

Summary of GreenHab operations: Additional information can be found in the GreenHab Report.
WATER USE: ~49 gallons (~5 gallons today)
Heater: On
Supplemental light: On
Harvest: 9 grams of microgreens and 21 grams of carrot tops (part of effort to thin carrot plot).

Summary of ScienceDome operations: N/A
Dual split: Per Mission Support’s message, the heat is on in the ScienceDome and set to 65 F.

Summary of RAM operations: No RAM operations today.

Summary of any observatory issues: No observatory issues.

Summary of health and safety issues: No health and safety issues.

Questions, concerns and requests to Mission Support: N/A

Journalist Report – December 7th

Sol: 10
Author: Dr. Kristine Ferrone, Commander
Title: Snow Day!

This morning the crew awoke to a winter wonderland outside; the landscape surrounding MDRS that had just started to become familiar was once again new and different with a soft white blanket over the ruddy rock formations. The crew consulted with Tier 1 mission support regarding their EVA plans for the day and decided that a short EVA to repair the mirror sample payload justified the additional risk of traversing the muddy terrain. Health and Safety Officer (Barbara Braun) and Crew Engineer (Ashley Kowalski) embarked in the late morning (with Green Hab Officer Matt Eby leading IVA comms) to the mirror payload location and worked through some realistic astronaut challenges such as working with tools and fasteners not specifically designed for use while wearing an EVA suit, bringing to mind the challenge Astronaut Mike Massimino experienced in removing tiny screws from a panel of the Hubble Space Telescope on a Space Shuttle EVA on STS-125. During the EVA, Technology Officer (Trevor Jahn) worked through challenges with data processing for Project Phantom while Commander (Kristine Ferrone) and XO (Alli Taylor) worked on planning, mission summary reports, and coordination for Friday’s STEM event. After the EVA, the crew settled in with hot chocolate to work on reports and begin preparation for end of mission activities. For dinner, Ashley and Trevor made extra large black bean burgers with microgreens Matt harvested from the Green Hab, served on homemade sourdough buns made by Barbara.

GreenHab Report – December 7th

GreenHab Officer: Matthew Eby
Environmental control: Heater on, door closed, 20% humidity.
Average temperatures: 68-deg F at time of watering. Range prior 24 hours: 64-deg F – 76-deg F
Hours of supplemental light: 4 hours
Daily water usage for crops: Approximately 5 gallons
Daily water usage for research and/or other purposes: 0 gallons
Water in Blue Tank: Approximately 195 gallons (Based on a 250 gallon capacity, visually 80% full)
Time(s) of watering for crops: 2:30 PM
Changes to crops: None
Narrative: Today plant food was added to the watering. Also, it was noted that the soil level was low on several of the tomato plants, exposing the top of the roots. More soil was added and additional supports provided to promote vertical growth (see pictures). Finally, a cage was added to one of the bean plants to provide support (see picture).
Harvest: 9 grams of microgreens and 21 grams of carrot tops (part of effort to thin carrot plot).
Support/supplies needed: None needed at this time.

Copyright © The Mars Society. All rights reserved. | Main Site